Rewrite without absolute value for the given conditions:
a)y=|x+5|, if x>−5
b)y=|x+5|, if x<−5
c)y=|x+5|, if x=−5
d)y=|x−3|+|x+2|−|x−5|, if x>5
e)y=|x−3|+|x+2|−|x−5|, if −2 f)y=|x−3|+|x+2|−|x−5|, if x<−2
g)y=|x−3|+|x+2|−|x−5|, if 3 h)y=|x−5|+|x+5|, if x>5
i)y=|x−5|+|x+5|, if x<−5
j)y=|x−5|+|x+5|, if −5 P.S if you solved all of these i will do anything you want :) Thanks again so much!

Answer by mathsolverplus(88)

For these questions, imagine the absolute value is not there and substitute any x-values within that domain. If it ends up being negative, you need to add a negative sign before the entire expression.
For example, for part a):
y=x+5 I am going to say that x = 0, which is bigger than -5, right?
y=5 The result is positive. Therefore, the expression does not
need a negative sign before it.
a) y=x+5
b) y=-(x+5) = -x-5
c) y=0
d) y=x-3+x+2-x-5 = x-6
e) y= -(x-3) + (x+2) + (x-5) = x
f) y= -(x-3) - (x+2) + (x-5) = -x-4
g) y= (x-3) + (x+2) + (x-5) = 3x-6
h) y= (x-5) + (x+5) = 2x
i) y= -(x-5) - (x+5) = -2x
j) y= -(x-5) + (x+5) = 10


For more additional instant FREE math help, please check out/follow our Instagram account !!!!!!
https://www.instagram.com/mathsolver.plus/